# Heisenberg limit in phase measurements: the threshold detection approach

### v1

@@ IV Two-arm interferometer
In this section, we adapt the analysis of the previous section to the antisymmetric two-arm interferometer shown in Fig. 2. Following Refs. [4, 23], we assume that the squeezed coherent state is injected into the first (bright) input port, and the squeezed vacuum state — into the second (dark) input port. This combinations gives the equal values of the mean optical power in two arms. We assume also that the interferometer is tuned in such a way that in the absence of the phase signal ($\phi=0$), both input states are reproduced at the respective bright and dark output ports (the dark fringe regime).
@@
$\hat{\mathcal{S}}_{1}=e^{R(\hat{a}_{1}^{\dagger}{}^{2}-\hat{a}_{1}^{2})/2}\,,$ (58a)
$\hat{\mathcal{S}}_{2}=e^{r(\hat{a}_{2}^{\dagger}{}^{2}-\hat{a}_{2}^{2})/2}\,,$ (58b)
are, respectively, the squeeze operators of the first and the second modes, and $\hat{a}_{1,2}$ and $\hat{a}_{1,2}^{\dagger}$ are the corresponding creation and annihilation operators of the input field (before the first beamsplitter). Similar to the single-arm case, we assume that the parameters $\alpha$, $R$ and $r$ are real (see also Ref. [3]).
We assume that the both beamsplitters are described by the following reflectivity/transmissivity matrix:
$\mathbb{B}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\,.$ (59)
In this case, the annihilation operators for the intracavity fields (before the phase shifts) are equal to
@@
For our calculations, we need the mean value of the sum photon number in the two arm
$N=\alpha^{2}+\mathop{\rm sinh}\nolimits^{2}R+\mathop{\rm sinh}\nolimits^{2}r\,,$ (61)
as well as the variance $(\Delta N_{-})^{2}$ of the differential photon number
It is easy to show that this variance is equal to
@@
Homodyne detector.
The input/output relations for the two-arm interferometer are the following, see again Ref. [3]:
$\begin{pmatrix}\hat{x}_{1\,\rm out}\\ \hat{p}_{2\,\rm out}\end{pmatrix}=\hat{\mathcal{U}}_{-}^{\dagger}(\phi)\begin{pmatrix}\hat{x}_{1}\\ \hat{p}_{2}\end{pmatrix}\hat{\mathcal{U}}_{-}(\phi)=\mathbb{R}(\phi)\begin{pmatrix}\sqrt{2}\alpha+\hat{x}_{1}e^{R}\\ \hat{p}_{2}e^{-r}\end{pmatrix},$ (64a)
$\begin{pmatrix}\hat{x}_{2\,\rm out}\\ \hat{p}_{1\,\rm out}\end{pmatrix}=\hat{\mathcal{U}}_{-}^{\dagger}(\phi)\begin{pmatrix}\hat{x}_{2}\\ \hat{p}_{1}\end{pmatrix}\hat{\mathcal{U}}_{-}(\phi)=\mathbb{R}(\phi)\begin{pmatrix}\hat{x}_{2}e^{r}\\ \hat{p}_{1}e^{-R}\end{pmatrix},$ (64b)
@@
is the two-mode evolution operator,
^$\hat{\mathcal{R}}_{-}(\phi)=e^{-i\hat{N}_{-}\phi}\,,$ (66)
is the two-mode phase shift operator, $\mathbb{R}(\phi)$ is again the phase rotation matrix (27), $\hat{x}_{1,2}$, $\hat{p}_{1,2}$ are the amplitude and phase quadrature operators of the input fields, defined by
@@
Similar to the single-arm scheme [compare Eqs. (24) and (64a)], we consider the homodyne measurement of the quadrature $\hat{p}_{2\,{\rm out}}$. In this case,
@@
where $(\Delta p_{2\,\rm out})^{2}$ is the variance of the measured quadrature and $G=\partial\langle{\hat{p}_{2\,\rm out}}\rangle/\partial\phi$ is the gain factor that again is equal to (31).
Let us minimize now Eq. (68) in $\alpha^{2}$ under the condition (61) at the best sensitivity point of $\phi=0$. In order to avoid over-cluttering the equations, we consider in more detail two most interesting particular cases: the canonical single-squeezed case of $R=0$, see Ref. [4], and the anti-symmetrically double-squeezed one of $R=-r$. The evident third option is the symmetrically double-squeezed one, $R=r$. However, calculations show that it provides the results similar to the considered in Sec. III single-arm case, but slightly inferior to it. Therefore, we do not include it in our consideration. The QCRB for this configuration was calculated in Ref. [24].
@@
In the second case, correspondingly,
@@
Threshold detector.
In the App. A, the analysis of the threshold detector, provided in Sec. III, is adapted to the two-mode case. In particular, it is shown there that the phase measurement error at $\phi=0$ is equal to
$(\Delta\phi_{0})^{2}=\frac{1}{4(\Delta N_{-})^{2}}\,.$ (75)
Let us minimize this value under the condition (61) for the same two particular cases that were considered for the homodyne measurement, namely $R=0$ and $R=-r$. If $R=0$, then the minimum is provided by
$\alpha^{2}=\frac{N+1/4}{2}+\mathcal{O}(N^{-1})$ (76)
@@
$(\Delta\phi_{0})^{2}=\frac{1}{4N(N+3/2)}+\mathcal{O}(N^{-4})\,.$ (77)
@@
The possible explicit form of $Y$ is developed is App. A using the same approach as in the previous (single-arm) case. The result is the following:
@@
Correspondingly, the value of this operator at the output of the interferometer is equal to
@@
Consider again the particular cases of $R=0$ and $R=-r$. In the first one, in order to simplify the equations, we assume that $|\phi|\ll 1$ and $N\gg 1$. The corresponding values of the gain factor $G$ and the variance of $Y_{\rm out}$ are calculated in App. A, see Eqs. (110). Substituting the optimization condition (76) into these equations, we obtain that
@@
In the second particular case of $R=-r$, the nonlinear term in Eq. (81) vanishes and $Y_{\rm out}$ becomes proportional to the quadrature $p_{2\,\rm out}$. This means that the homodyne measurement of this quadrature is now the optimal procedure, described by the phase measurement error (74).
V Possible implementation of the threshold detection
The goal of this section is to show that the measurement of $\hat{Y}$ in principle can be implemented using the quadratic ($\chi^{(2)}$) optical nonlinearity. The correponding scheme is shown in Fig. 3.
Figure 3: A possible implementation of the threshold detector. HD: homodyne detector, BS: asymmetric beamsplitter with the amplitude transmissivity $T\to 0$; $|{\beta/T}\rangle$: coherent state with the amplitude $\beta/T$.
Omitting for brevity the subscripts “out”, the operator (52) can be presented as follows:
$\hat{Y}=A\hat{p}+Bx\circ p$ (83)
where $A$ and $B$ are the real $c$-numbers. Apply to $\hat{Y}$ the displacement operator $\hat{\mathcal{D}}(\beta)$, where
$\beta=-\frac{A}{\sqrt{2}B}\,.$ (84)
This linear transformation can be easily implemented by mixing the measured field with the strong coherent one, using a non-symmetric beamsplitter (BS in Fig. (3)) with the amplitude transmissivity $T\to 0$. Taking into account that
$\hat{\mathcal{D}}^{\dagger}(\beta)\hat{x}\hat{\mathcal{D}}(\beta)=\hat{x}+\beta\sqrt{2}\,,\quad\hat{\mathcal{D}}^{\dagger}(\beta)\hat{p}\hat{\mathcal{D}}(\beta)=\hat{p}\,,$ (85)
we obtain:
$\hat{\mathcal{D}}^{\dagger}(\beta)\hat{Y}\hat{\mathcal{D}}(\beta)=B\hat{x}\circ\hat{p}=\frac{B}{2i}(a^{2}-a^{{\dagger}2})\,.$ (86)
Below, we omit for brevity the factor $B$ and consider measurement of the operator
$\hat{M}=\frac{1}{2i}(a^{2}-a^{{\dagger}2})$ (87)
Suppose that the output light of the interferometer interacts with another (probe) mode by means of the second order optical nonlinearity, see the block $\chi^{(2)}$ in Fig. 3. This interaction is described (in the rotating wave picture) by the Hamiltonian
$\hat{H}=\hbar g(\hat{b}^{\dagger}\hat{a}^{2}+\hat{a}^{{\dagger}2}\hat{b})=\frac{\hbar g}{\sqrt{2}}\bigl{[}(\hat{a}^{2}+\hat{a}^{{\dagger}2})\hat{X}-i(\hat{a}^{2}-\hat{a}^{{\dagger}2})\hat{P}\bigr{]},$ (88)
where
$\hat{b}=\frac{\hat{X}+i\hat{P}}{\sqrt{2}}$ (89)
is the annihilation operator of the probe mode. The corresponding Heisenberg equation of motion for $\hat{X}$ is the following:
$\dfrac{d\hat{X}}{dt}=\sqrt{2}g\hat{M}\,.$ (90)
Now suppose that the interaction time $\tau$ is short:
$g\tau\ll 1\,.$ (91)
In this case, keeping only linear in $\tau$ terms in the solution to this equation, we obtain:
$\hat{X}(\tau)=\hat{X}+\sqrt{2}g\tau\hat{M}\,.$ (92)
where $\hat{X}(\tau)$ is the value of $X$ after the interaction. Therefore, measurement of $\hat{X}(\tau)$ allows to obtain the information on $\hat{M}$ with the precision, depending on the initial uncertainty $\Delta X$ of $X$. In principle, preparing the probe mode in a highly squeezed state with $\Delta X\to 0$, it is possible to measure $\hat{M}$ and therefore $\hat{Y}$ with arbitrary high precision.
VI Discussion
@@
Two-arm, $R=0$ $\dfrac{1}{4N(N+1)}$ $\dfrac{1}{2N+1}$ $\approx\dfrac{1}{4N(N+3/2)}$ $\approx\dfrac{2}{9N+5/4}$
Two-arm, $R=-r$ $\dfrac{1}{2N(N+2)}$ $\sim 1$ $\dfrac{1}{2N(N+2)}$ $\sim 1$
@@
Table 1: Optimized sensitivity for the three considered interferometric configurations and for the homodine detection (HD) and threshold detection (TD) measurement schemes.
@@
In the most important from the practical point of view case of the small signal phase shift $|\phi|\ll 1$, the sensitivity limits (33), (54), (71), (74), and (82), calculated in Sections III and IV, can be presented in the following unified form:
$(\Delta\phi)^{2}\approx(\Delta\phi_{0})^{2}\times\biggl{(}1+\frac{\phi^{2}}{(\delta\phi)^{2}}\biggr{)},$ (93)
where $\Delta\phi_{0}$ is the value of the measurement error $\Delta\phi$ at the best sensitivity point of $\phi=0$ and $\delta\phi$ is the half-width at half minimum (HWHM) of the range as of values of $\phi$ within which the measurement error does not exceed $\sqrt{2}\,\Delta\phi_{0}$.
The factor $\delta\phi$ allows to some extent to take into account the a priori information on the phase. If the a priori distribution fits into the range $|\phi|<\delta\phi$, then the resulting measurement error can be approximated by $\Delta\phi$. In the opposite case, however, the more sophisticated Bayesian or maximum likelood approaches should be used.
The corresponding values of $\Delta\phi_{0}$ and $\delta\phi$ are presented in Table 1. It can be seen that in all cases, the HL scaling can be achieved, but with different numerical factors. The best sensitivity at $\phi=0$ is provided by the single-arm interferometer combined with the threshold detection. At the same time, the single-arm interferometer is characterized by the most narrow range of the good sensitivity, $\delta\phi\sim 1/N$, both for the homodyne and threshold detection cases.
@@
Actually, these two features stem from the same origin, namely the strongly enhanced by squeezing amplitude quadrature of the input light $\hat{x}e^{r}$, which appears in the phase quadrature of the output light $\hat{p}_{\rm out}$ if $\phi\neq 0$, see Eq. (24). This leads to the sharp increase of the variance $\Delta p_{\rm out}$ with the increase of $\phi$, limiting the value of $\delta\phi$, see Eq. (30). At the same time, this means also that $\Delta p_{\rm out}$ carries information on $\phi$ that can be retrieved by some non-linear measurement, in particular, by the threshold detector. It is interesting that the optimization procedure prefers this information to the linear one, setting the value of the classical amplitude $\alpha$ equal to zero, see Eq. (42). As a result, the threshold detector provides smaller by $\sqrt{2}$ value of $\Delta\phi_{0}$ in comparison with the linear homodyne detector.
In the single-squeezed ($R=0$) two-arm interferometer case, the amplitude quadrature $\hat{x}_{1}$ of the input light of the first mode appears in the measured quadrature $\hat{p}_{\rm 2\,out}$, see Eq. (64a). In principle, the above consideration is valid in this case too. However, the variance of $\hat{x}_{1}$ now is much smaller — it corresponds to the vacuum state, see Eq. (69). Therefore, both discussed above effects are much less pronounced. The good sensitivity range now scales as $\delta\phi\sim 1/\sqrt{N}$, but the threshold detector provides only marginal advantage in the sensitivity in comparison with the homodyne one.
@@
Finally, in the double-squeezed case of $R=-r$, the variance of $\hat{p}_{\rm 2\,out}$ does not depend on $\phi$ at all, see Eq. (72). As a result, the non-linear term in $\hat{Y}_{\rm out}$ vanishes and the threshold detection scheme reduces to the ordinary homodyne measurement. The good sensitivity range in this case is the broadest one, $\delta\phi\sim 1$. At the same time, the value of $\Delta\phi_{0}$ is $\sqrt{2}$ times bigger than in the previous case. This feature can be attributed to the additional term $\mathop{\rm sinh}\nolimits^{2}R=\mathop{\rm sinh}\nolimits^{2}r$ in the expression (61) for the mean number of quanta that leads to the smaller value of $\alpha$ for the same values of $N$ and $r$. It is worth noting that the corresponding HL prefactor $1/2$ is still smaller than $\approx\!\!1.38^{2}$ of Refs. [15, 16, 17], despite the exotic non-Gaussian states considered in that works.
Appendix A Threshold detection in the two-arm interferometer case

### v2

@@ IV Two-arm interferometer
The antisymmetric two-arm interferometer which we consider here is shown in Fig. 2. We assume that both beamsplitters are described by the following reflectivity/transmissivity matrix:
$\mathbb{B}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, (41)
and in the absence of the phase shifts $\pm\phi$, the arm lengths are equal to each other. In this case, at $\phi=0$ both input states are reproduced at the respective bright and dark output ports (the dark fringe regime).
Note that in literature, the signal phase shifts are often denoted as $\pm\phi/2$. Here we use the $\pm\phi$” convention, because it is more consistent with the single arm case, providing the same values of the SNL (1) and the squeezing-enhanced SNL (2) as the single-arm interferometer.
Following Refs. [4, 33], we assume that the squeezed coherent state is injected into the first (bright) input port, and the squeezed vacuum state — into the second (dark) input port. This combination gives the equal values of the mean optical power in two arms, keeping the (anti-) symmetry of the scheme intact, see Refs.  [34, 3].
@@
$\hat{\mathcal{S}}_{1}=e^{R(\hat{a}_{1}^{{\dagger}\,2}-\hat{a}_{1}^{2})/2}\,,\quad\hat{S}_{2}=e^{r(\hat{a}_{2}^{{\dagger}\,2}-\hat{a}_{2}^{2})/2}$ (44)
are, respectively, the squeeze operators of the first and second modes, and $\hat{a}_{1,2}$ and $\hat{a}^{\dagger}_{1,2}$ are the corresponding creation and annihilation operators of the input field (before the first beamsplitter). Similarly to the single-arm case, we assume that the parameters $\alpha$, $R$ and $r$ are real (see also Ref. [3]). It is easy to show that in this case, the mean value of the sum photon number in the interferometer is equal to
$N=\alpha^{2}+\mathop{\rm sinh}\nolimits^{2}R+\mathop{\rm sinh}\nolimits^{2}r\,.$ (45)
The phase shift operator (9) in the antisymmetric two-arm case has the following form:
$\hat{\mathcal{R}}_{-}(\phi)=e^{-i\hat{N}_{-}\phi}\,$, (46)
where
is the difference of the photon numbers in the arms and
@@
are the annihilation operators for the intracavity fields before the phase shifts, see Eq. (41). It follows from Eqs. (42) and (47) that the variance of $\hat{N}_{-}$ is equal to
@@
In the Heisenberg evolution picture, the input/output relations for the two-arm interferometer are the following, see Ref. [3]:
$\begin{pmatrix}\hat{x}_{1\,\rm out}\\ \hat{p}_{1\,\rm out}\\ \hat{x}_{2\,\rm out}\\ \hat{p}_{2\,\rm out}\end{pmatrix}=\hat{\mathcal{U}}_{-}^{\dagger}(\phi)\begin{pmatrix}\hat{x}_{1}\\ \hat{p}_{1}\\ \hat{x}_{2}\\ \hat{p}_{2}\end{pmatrix}\hat{\mathcal{U}}_{-}(\phi)=\begin{pmatrix}(\sqrt{2}\alpha+\hat{x}_{1}e^{R})\cos\phi+\hat{p}_{2}e^{-r}\sin\phi\\ -\hat{x}_{2}e^{r}\sin\phi+\hat{p}_{1}e^{-R}\cos\phi\\ \hat{x}_{2}e^{r}\cos\phi+\hat{p}_{1}e^{-R}\sin\phi\\ -(\sqrt{2}\alpha+\hat{x}_{1}e^{R})\sin\phi+\hat{p}_{2}e^{-r}\cos\phi\end{pmatrix}$, (50)
@@
is the two-mode evolution operator, $\hat{x}_{1,2}$, $\hat{p}_{1,2}$ are the amplitude and phase quadrature operators of the input fields, defined by
@@
Homodyne detector.
Similarly to the single-arm scheme [compare Eqs. (24) and (50)], we consider the homodyne measurement of the quadrature $\hat{p}_{2\,{\rm out}}$ which carries the major part of the phase information, see Eq. (50). In this case,
@@
where
$(\Delta p_{2\,\rm out})^{2}=\frac{1}{2}\bigl(e^{-2r}\cos^{2}\phi+e^{2R}\sin^{2}\phi\bigr)$ (54)
is the variance of the measured quadrature and $G=\partial\langle{\hat{p}_{2\,\rm out}}\rangle/\partial\phi$ is the gain factor that again is equal to (30).
Note that now, opposite to the single-arm case, two components of $\Delta p_{2\,\rm out}$ originate from two independent input modes and does not bound by the uncertainty relation. Therefore, increase of the squeeze factor $r$ does not affect the second term in Eq. (54), proportional to $e^{2R}$. As a result, much broader range $\delta\phi$ could be achieved in the two-arm case.
Let us minimize Eq. (53) in $\alpha^{2}$ under the condition (45) at the best sensitivity point of $\phi=0$. In order to avoid over-cluttering of the equations, we consider in more detail two most interesting particular cases: the canonical single-squeezed case of $R=0$ and the anti-symmetrically double-squeezed one of R=rR=-r. The evident third option is the symmetrically double-squeezed one, $R=r$. However, calculations show that it provides the results similar to the considered in Sec. III single-arm case, but slightly inferior to it. Therefore, we do not consider it here.
@@
In the second case, $\Delta p_{2\,\rm out}$ does not depend on $\phi$:
@@
QCRB and threshold detector at $\phi\to 0$.
It follows from Eqs. (9) and (46) that the QCRB in the double-arm interferometer case is equal to
$(\Delta\phi)^{2}=\frac{1}{4(\Delta N_{-})^{2}}\,$, (61)
where the photon number variance is given by Eq. (49).
We minimize this value under the condition (45) for the same two particular cases that were considered for the homodyne measurement, namely $R=0$ and R=rR=-r. If $R=0$, then the minimum is provided by
$\alpha^{2}\approx\frac{N+1/4}{2}$ (62)
@@
$(\Delta\phi_{0})^{2}\approx\frac{1}{4N(N+3/2)}\,.$ (63)
@@
Threshold detector at $\phi\neq 0$.
In order to calculate the measurement error for $\phi\neq 0$, we again need the explicit form of the operator $\hat{Y}$. It is shown in App. B that it has the following, again bilinear, form:
@@
Consider again the same two particular cases of $R=0$ and R=rR=-r. In the first one, in order to simplify the equations, we assume that $|\phi|\ll 1$. The corresponding values of the gain factor $G$ and the variance of $Y_{\rm out}$ are calculated in App. B, see Eqs. (96). Substituting the optimization condition (62) into these equations, we obtain that
@@
In the second particular case of R=rR=-r, the nonlinear term in Eq. (67) vanishes and $Y_{\rm out}$ becomes proportional to the quadrature $p_{2\,\rm out}$. This means that the homodyne measurement of this quadrature is now the optimal procedure, described by the phase measurement error (60).
V Discussion
@@
Two-arm, $R=0$ $\dfrac{1}{4N(N+1)}$ $\dfrac{2}{N+\nicefrac{{1}}{{2}}}$ $\approx\!\!\dfrac{1}{4N(N+\nicefrac{{3}}{{2}})}$ $\approx\!\!\dfrac{8}{9N+\nicefrac{{5}}{{4}}}$
Two-arm, R=rR=-r $\dfrac{1}{2N(N+2)}$ $\sim 1$ $\dfrac{1}{2N(N+2)}$ $\sim 1$
@@
Table 1: Optimized sensitivity for the three considered interferometric configurations and for the homodyne detection (HD) and threshold detection (TD) measurement schemes; $m$: the configuration number.
@@
In the most important from the practical point of view case of the big mean photon number $N\gg 1$ and the small signal phase shift $|\phi|\ll 1$, the sensitivity limits (32), (39), (57), (60), and (68), calculated in Sections III and IV, can be presented in the following unified form:
$(\Delta\phi)^{2}\Bigr|_{\phi}\approx(\Delta\phi_{0})^{2}\times\biggl(1+\frac{4\phi^{2}}{(\delta\phi)^{2}}\biggr).$ (69)
The values of $\Delta\phi_{0}$ and $\delta\phi$ for the configurations considered in this paper are presented in Table 1. It can be seen that in all cases, the HL scaling (3) can be achieved, but with different values of the factor $K$. It can be seen also that there is an interdependence between values the values of $\Delta\phi_{0}$ and $\delta\phi$. In all cases except for the single-arm interferometer with the homodyne detection, it can be roughly estimated as follows:
$\log_{N}(\delta\phi)\sim\log_{2}(N\Delta\phi_{0})+\frac{1}{2}\,.$ (70)
In almost all considered cases, the ordinary homodyne detector provides virtually the same sensitivity, both in terms of $\Delta\phi_{0}$ and $\delta\phi$, as the non-linear threshold detector. The only exception is the single arm interferometer, where threshold detector gives twice as small value of $(\Delta\phi_{0})^{2}$, which corresponds to the best sensitivity among all considered here cases. At the same time, the single-arm interferometer is characterized by the most narrow range of high sensitivity, $\delta\phi\sim 1/N$ for both the homodyne and threshold detection cases. Note that it is close to $\Delta\phi_{0}$, that is, in this case the a priori knowledge about the phase distribution before the measurement should be as good as measurement precision.
@@
These two features stem from the same origin, namely the enhanced by the squeezing amplitude quadrature of the input light $\hat{x}e^{r}$, which appears in the measured phase quadrature of the output light $\hat{p}_{\rm out}$ at $\phi\neq 0$, see Eq. (24). Note also that this term is bound by the uncertainty relation with the term $\hat{p}e^{-r}$ which defines the measurement error at $\phi=0$. As a result, the variance $\Delta p_{\rm out}$ sharply increases with the increase of $\phi$, limiting the value of $\delta\phi$.
At the same time, the dependence of $\Delta p_{\rm out}$ on $\phi$ can be exploited as the additional source of information using more advanced data processing or more sophisticated measurements. In particular, in Ref. [24] the use of the maximum likelihood method was considered. The disadvantage of this approach is that it requires multiple measurements, which leads to the additional factor $\sqrt{p}$ in the numerator of the HL (3), where $p$ is the repetitions number. Another approach is the use of a non-linear measurement, in particular, of the threshold detector. It is interesting that in this case the optimization procedure prefers the variance information to the linear one, setting the value of $\alpha$ equal to zero, see Eq. (34).
In the two-arm interferometer case, the measured output quadrature $\hat{p}_{\rm out}$ depends on the quadratures $\hat{x}_{1}$ and $\hat{p}_{2}$ of the two independently prepared input modes and does not bound by the uncertainty relation. As a result, the squeezing of $\hat{p}_{2}$ does not affect uncertainty of $\hat{x}_{1}$, and vice versa. In particular, if the first mode is not squeezed, $R=0$, then the value of $\delta\phi$ scales only as $1/\sqrt{N}$. This feature allows to use, before the main measurement, a preliminary “ranging” shot noise limited one (without squeezing in both arms), which is characterized by $\delta\phi\sim 1$, thus providing, in the series of two measurement, both broad $\delta\phi$ and the HL sensitivity.
@@
Finally, in the double-squeezed case of R=rR=-r, the variance of $\hat{p}_{\rm 2\,out}$ does not depend on $\phi$ at all, see Eq. (58). As a result, the non-linear term in $\hat{Y}_{\rm out}$ vanishes and the threshold detection scheme reduces to the ordinary homodyne measurement. The high sensitivity range in this case is the broadest one, $\delta\phi\sim 1$. At the same time, the value of $(\Delta\phi_{0})^{2}$ is twice as big as in the previous case. This feature can be attributed to the additional term $\mathop{\rm sinh}\nolimits^{2}R=\mathop{\rm sinh}\nolimits^{2}r$ in the expression (45) for the mean number of quanta that leads to the smaller value of $\alpha$ for the same values of $N$ and $r$. However, it is worth noting that the corresponding HL prefactor $1/2$ is still smaller than $\approx\!\!1.38^{2}$ of the limit (4), despite the exotic non-Gaussian nature of the quantum states providing that limit.
Acknowledgements.
This work of D.S., V.G., B.N., and F.K. was supported by the Theoretical Physics and Mathematics Advancement Foundation “BASIS” Grant #23-1-1-39-1. Authors thank L. Pezze for his valuable remarks on this work.
Appendix A Threshold detection, single-arm interferometer
Threshold detector at $\phi=0$.
In the Schrödinger picture, the output field of the interferometer is described by the following density operator:
where
$\hat{\rho}_{0}=|{\psi_{0}}\rangle\langle{\psi_{0}}|\,.$ (72)
By substituting it into Eq. (14), we obtain:
$i[\hat{\rho}_{0},\hat{N}]=\hat{\rho}_{0}\circ\hat{Y}\,$. (73)
Consider then a set of quantum states $\{|{\psi_{k}}\rangle\}$, with $k=1,2,\dots$, complementing $|{\psi_{0}}\rangle$ to the full orthonormal set:
In this representation, Eq. (73) can be rewritten as follows:
This equation can be solved explicitly, giving:
$\langle{\psi_{k}}|\hat{Y}|{\psi_{l}}\rangle=\begin{cases}0\,,&k=l=0\,,\\ 2i\langle{\psi_{0}}|\hat{N}|{\psi_{l}}\rangle\,,&k=0\,,\ l>0\,,\\ -2i\langle{\psi_{k}}|\hat{N}|{\psi_{0}}\rangle\,,&k>0\,,\ l=0\,,\\ \text{undefined}\,,&k>0\,,\ l>0\,.\end{cases}$ (76)
It is easy to see from this equation, that
$(\Delta Y)^{2}=4(\Delta N)^{2}$ (77)
With an account for Eq. (16), we obtain Eq. (33).
Threshold detector at $\phi\neq 0$.
In order to calculate the measurement error for other values of $\phi$, the explicit form of the operator $\hat{Y}$ is required, that is, the set $\{|{\psi_{k}}\rangle\}$ and the fourth clause in Eq. (76) have to be defined explicitly. In this paper, following Refs. [31, 32, 29], we consider a simple bilinear in $\hat{x}$, $\hat{p}$ form of $\hat{Y}$ that, in principle, could be implemented in an experiment, see Ref. [30].
where the operators $\hat{\mathcal{D}}$ and $\hat{\mathcal{S}}$ are given by Eqs. (19) and (20), respectively. In this case, we obtain that if $l>0$, then
where
As a result, Eq. (76) takes the following form:
[…]
The form of $\hat{Y}$ that satisfies Eq. (81) for all $k>0,l>0$ is the following:
$\hat{\mathcal{S}}^{\dagger}\hat{\mathcal{D}}^{\dagger}\hat{Y}\hat{\mathcal{D}}\hat{\mathcal{S}}=\hat{\mathcal{Y}}+\hat{\mathcal{Y}}^{\dagger}+\hat{a}^{\dagger}\hat{Q}\hat{a}=-2\hat{p}\circ(\sqrt{2}\alpha e^{r}+\hat{x}\mathop{\rm sinh}\nolimits 2r)+\hat{a}^{\dagger}\hat{Q}\hat{a}\,$, (82)
where $\hat{Q}$ is an arbitrary Hermitian operator.
The last term in Eq. (82) arises due to the non-uniqueness of the operator $\hat{Y}$, briefly discussed in Ref. [29]. The choice of $\hat{Q}$ does not affect the best sensitivity value $\Delta\phi_{0}$, defined by the QCRB (10), but could affect the values of $\Delta\phi$ for $\phi\neq 0$. In this paper, we consider the case of $\hat{Q}=0$, which corresponds to simplest bilinear form of $\hat{Y}$.
Rolling back the squeeze and displacement operators, we obtain Eq. (36).
Appendix B Threshold detection, two-arm interferometer case